OpenStudy (anonymous):

Point P is located at (−2, 7) and point R is located at (1, 0). Find the y value for the point Q that is located the distance from point P to point R. 4.9 4.7 2.5 2.3 notice me senpai

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

:)

OpenStudy (compassionate):

Point P is located at (−2, 7) and point R is located at (1, 0). Find the y value for the point Q that is located the distance from point P to point R. _________________________________________ [(-2, 7), (1, 0)] d = 2/3 Basically, find the difference between your x and y terms. The difference of -2 and 1 is 3, The difference of 7 and 0 is 7. x = 3 y = 7 Multiply it by two thirds. 2/3 * 3 = 2 2/3 * 7 = 4.7 (;

OpenStudy (anonymous):

This is the wrong! The answer is 2.3, the key thing to look at is that the questions is asking "from point P to R", not "R to P" (which 4.7 is point R to P). Be careful when answering these questions.